OpenStudy (anonymous):

a=b+c-2bc x cos A ? Simplify.. trig?

OpenStudy (alexwee123):

solve for A?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

that is a^2=b^2+c^2-2bc c0s A

OpenStudy (anonymous):

Explain pleasee

OpenStudy (anonymous):

law of cosines right?

OpenStudy (anonymous):

and it cannot mean "simplify" maybe it means "solve for \(A\)"

OpenStudy (anonymous):

\[a^2=b^2+c^2-2bc\cos(A)\] \[a^2-b^2-c^2=-2bc\cos(A)\] \[\frac{b^2+c^2-a^2}{2bc}=\cos(A)\] \[A=\cos^{-1}\left(\frac{b^2+c^2-a^2}{2bc}\right)\]

OpenStudy (anonymous):

OMG thanks so much!!

OpenStudy (anonymous):

yw